# Reservoir-assisted quantum battery charging at finite temperatures

### v2

@@ V Summary
In this study, we investigate charging models for single-particle and multi-particle quantum batteries based on homodyne quantum feedback control. In the single-particle model, both the charger and the battery are modeled as two-level atomic systems. Quantum feedback control is applied to the charger during the charging process, enabling energy transfer through the coupling between the charger and the battery. Subsequently, we examined the effects of bosonic and fermionic thermal reservoirs on the performance of the single-particle quantum battery. The findings reveal the existence of optimal charging parameters for the system. Compared to bosonic thermal reservoirs, fermionic reservoirs at finite temperature facilitate energy storage and extraction more efficiently. Moreover, the battery charging efficiency $R$ improves with increasing temperature. In the multi-particle quantum battery model, the charger is still modeled as a two-level system, while the quantum battery consists of an ensemble of two-level systems. We further analyze the charging process when the multi-particle quantum battery is collectively placed in either a bosonic or fermionic thermal reservoir. Additionally, we investigate the effects of the reservoir temperature $T$, the interparticle coupling strength $J$, and the number of particles $N$ on the performance of the multi-particle quantum battery. The results indicate that when the multi-particle quantum battery is placed in a bosonic thermal reservoir, increasing the temperature reduces both the energy density and the average ergotropy, consistent with the behavior observed in the single-particle quantum battery. Furthermore, expanding the size of the quantum battery in a bosonic thermal reservoir requires satisfying three conditions: a lower thermal reservoir temperature, weak inter-atomic interactions within the battery, and a low dissipation rate of the quantum battery. For the multi-particle quantum battery in a fermionic thermal reservoir, high temperatures still exhibit corresponding advantages in terms of the battery’s energy density and average ergotropy. Simultaneously, the conditions for constructing a multi-particle quantum battery are significantly relaxed. Even under strong interparticle interactions and higher battery dissipation rates, the multi-particle quantum battery still maintains relatively high-performance levels in both energy storage and energy extraction under high-temperature conditions.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12174048; Shenyang Ligong University 2024 introduced high-level talent research support funds under Grant No. 1010147001309.
@@
Then setting $\dot{\rho}(t)=0$, we obtained the following steady-state solutions
@@
$\mathcal{E}(t)$ $=$ $\frac{\omega_{0}}{2}\{\sqrt{4|\rho_{12}(t)+\rho_{34}(t)|^{2}+[2(\rho_{11}(t)+\rho_{33}(t))-1]^{2}}$ (44)
$+[2(\rho_{11}(t)+\rho_{33}(t))-1]\}.$
According to Eqs. (43) and (44), the energy storage and the ergotropy of the quantum battery in the steady state can be determined. The expression for the ergotropy of the battery is more complex and will not be provided here, while the expression for energy storage is as follows
@@
When the quantum battery is in contact with a fermionic thermal reservoir, the differential equations are similarly obtained by solving the corresponding master equation. By setting $\dot{\rho}(t)=0$, the steady-state solution is obtained as follows
@@
We set $\delta=f/\Gamma_{C}$ for simplicity. The other steady-state expansion coefficients are all zero, i.e. $\rho_{12}^{\infty}=\rho_{21}^{\infty}=\rho_{13}^{\infty}=\rho_{31}^{\infty}=\rho_{24}^{\infty}=\rho_{42}^{\infty}=\rho_{34}^{\infty}=\rho_{43}^{\infty}=0$. According to Eqs. (43) and (44), the stored energy and the ergotropy of the quantum battery in steady state can be determined. The expression for the ergotropy of the battery is more complex and will not be provided here, while the expression for energy storage is as follows
@@
Figure 22: The steady-state stored energy (solid line) and the ergotropy (dashed line) of the quantum battery are plotted as functions of the fermionic thermal reservoir temperature under optimal charging parameters of $\Gamma_{C}=2g$ and $\delta=f/\Gamma_{C}=1$. The coupling strength between the charger and the battery is set to $g=0.01\omega_{0}$. The red curve represents the results from our proposed scheme. The corresponding results from Ref. Song et al. (2024b) are shown as the blue curve in the inset, with the parameter choices matching those in Fig. 5(a) of Ref. Song et al. (2024b), i.e.,$F=\Gamma_{C}=4g$.
Figure 23: The multi-particle quantum battery is placed in a bosonic thermal reservoir, with the strong dissipation rate of $\Gamma_{B}=0.5g$. The optimal energy density of the battery is plotted as a function of the number of quantum battery particles, under different reservoir temperatures and internal interactions between particles within the battery. The remaining information is provided in Fig. 14.
@@
Figure 25: The multi-particle quantum battery is in a bosonic thermal reservoir. The charging efficiency $R$ of the battery is shown as a function of the number of battery particles $N$. The remaining information is the same as in Fig. 16.
Appendix B A comparative analysis between the current work and reference Song et al. (2024b)
In reference Song et al. (2024b), the charger is modeled as a dissipative system embedded in either a bosonic or fermionic thermal reservoir, and powered by a driving laser field, while the quantum battery is isolated. For the case where the charger is embedded in a fermionic thermal reservoir with positive temperatures, we evaluate the quantum battery’s performance and conduct a comparative analysis with our proposed scheme. When the system reaches a steady state, the stored energy $E_{B}(\infty)$ of the battery and the ergotropy $\mathcal{E}(\infty)$ of the battery are,
$\mathcal{E}(\infty)$ $=$ $\frac{\omega_{0}}{2}\big{(}-1+2\sqrt{\frac{g^{2}[B+g^{2}(1-4n_{f})](1-2n_{f})^{2}\Gamma_{C}^{4}(2g^{2}+\Gamma_{C}^{2})^{2}}{\{4F^{4}g^{2}+2[A+2g^{4}(1-2n_{f})]\Gamma_{C}^{2}+[B+2g^{2}(1-2n_{f})]\Gamma_{C}^{4}\}^{2}}}$ (53)
$+\frac{4F^{4}g^{2}+2A\Gamma_{C}^{2}+B\Gamma_{C}^{4}}{8F^{4}g^{2}+4[A+2g^{4}(1-2n_{f})]\Gamma_{C}^{2}+2[B+2g^{2}(1-2n_{f})]\Gamma_{C}^{4}}\big{)},$
@@
where $A=F^{2}g^{2}[1-8(n_{f}-1)n_{f}]+F^{4}(1-2n_{f})^{2}+4g^{4}n_{f}$, and $B=F^{2}(1-2n_{f})^{2}+4g^{2}n_{f}$. Here, $F$ is the amplitude of the driving field, $\Gamma_{C}$ is the dissipation rate of the charger, $g$ is the coupling strength between the charger and the battery, and $n_{f}$ is the average excitation number of the fermionic thermal reservoir.
@@
Figure 26: The multi-particle quantum battery is placed in a fermionic thermal reservoir. The optimal energy density of multi-particle quantum batteries is shown as a function of the particle number $N$. The only difference from Fig. 18 is that the dissipation rate of the battery is set to $\Gamma_{B}=0.5g$.
[…]
@@
blueTo better compare the differences between our scheme and that in the Ref. Song et al. (2024b), Fig. 22 plots the steady-state stored energy of the quantum battery (red solid line) and its ergotropy (red dashed line) under our charging scheme in a fermionic thermal reservoir as a function of temperature. The corresponding results from Ref. Song et al. (2024b) are displayed in the inset, with the parameter choices matching those in Fig. 5(a) of Ref. Song et al. (2024b), i.e., $F=4g$ and $\Gamma_{C}=4g$. Here, the solid line represents the steady-state energy storage of the battery, and the dashed line corresponds to the maximum extractable energy (the ergotropy) of the battery. The results indicate that, in our scheme, both the energy storage and the maximum extractable energy of the battery increase with the temperature rise. In contrast, the results in Ref. Song et al. (2024b) show that, as the temperature increases, the maximum stored energy $E_{B}^{\rm optimal}(\infty)$ of the quantum battery tends to approach $0.5\omega_{0}$, and the maximum extractable energy $\mathcal{E}(\infty)$ gradually decreases to zero.
@@
Appendix C A strongly dissipative multi-particle quantum battery in a bosonic thermal reservoir
@@
Figure 27: In a fermionic thermal reservoir, the average ergotropy of the multi-particle quantum battery in the steady state is shown as a function of the particle number $N$. Except for $\Gamma_{B}=0.5g$, all other parameters are the same as in Fig. 19.
@@
Figure 28: The multi-particle quantum battery is coupled to a fermionic thermal reservoir. The charging efficiency $R$ of the battery is shown as a function of the number of battery particles $N$, where $\Gamma_{B}=0.5g$, and the other parameters are the same as those in Fig. 20.
@@
In this section, we investigate the dependence of the energy density, the average ergotropy, and the charging efficiency $R$ of the multi-particle quantum battery on three key parameters: the thermal reservoir temperature, the number of particles $N$ and the interatomic interaction strength $J$, under strong dissipation conditions ($\Gamma_{B}=0.5g$) in a bosonic thermal reservoir. Figure 23 shows how the energy density of the multi-particle quantum battery varies with the particle number $N$ under different temperatures and interatomic interactions within the battery. Panels (a)-(c) correspond to $J=0$, $J=0.1g$, and $J=g$, respectively, and the different curves represent different thermal reservoir temperatures. Figures 23 (d)-(f) correspond to the optimal dissipative rates of the charger required for Figs. 23 (a)-(c). Unlike the weak dissipation regime depicted in Fig. 14, the strong dissipation regime exhibits a universal decrease in energy density with increasing particle number $N$, independent of both the thermal reservoir temperature (ranging from low to high) and the strength of interparticle coupling within the battery (whether weak or strong). Consequently, constructing multi-particle quantum batteries in a strongly dissipative bosonic thermal reservoir is not advisable. However, it should be noted that for the number of particles $N>2$, a high-temperature environment can alleviate the reduction in energy density induced by increasing $N$. Figure 24 shows correspondingly the behavior of the average ergotropy of the multi-particle quantum battery. The result shows that in a high-temperature environment, an increase in the number of particles $N$ is beneficial for energy extraction. Meanwhile, as shown in Fig. 25, the charging efficiency $R$ of the battery increases with the number of particles under high-temperature conditions. In contrast, under low-temperature conditions, $R$ decreases as the number of particles $N$ increases.
@@
Appendix D A strongly dissipative multi-particle quantum battery in a fermionic thermal reservoir.
@@
When the multi-particle quantum battery operates in a strongly dissipative fermionic thermal reservoir ($\Gamma_{B}=0.5g$), the energy density of the battery as a function of the number of particles $N$ is shown in Fig. 26. Figures 26 (a)-(c) correspond to $J=0$, $J=0.1g$, and $J=g$, respectively, with different curves representing various temperatures of the thermal reservoir. Figures 26 (d)-(f) show the corresponding optimal dissipative rates of the charger required for the cases in Fig. 26 (a)-(c), respectively. The results indicate that in a strong fermionic thermal reservoir, an increase in temperature enhances the energy density of the quantum battery. This means that higher temperatures facilitate energy storage in multi-particle quantum batteries. However, increasing the number of particles in the battery under low temperature conditions will have a negative impact on the energy density, regardless of whether the interatomic interaction strength is strong or weak. Conversely, under high-temperature conditions, as the number of battery particles increases, the battery will maintain a state with higher energy density, and the interatomic interaction strength $J$ has a relatively minor influence on its energy density. As shown in Fig. 27, the behavior of the average ergotropy closely resembles that of energy storage. Furthermore, Fig. 28 shows that the charging efficiency $R$ of the battery increases with the number of particles under high-temperature conditions, whereas under low-temperature conditions, $R$ decreases as $N$ increases.
@@
Overall, regardless of whether the multi-particle quantum battery operates under a low or high dissipation rate, its performance in a high-temperature fermionic thermal reservoir is significantly superior to that of a counterpart constructed in a bosonic thermal reservoir.

### v3

@@ V Summary
In this study, we investigated charging models for single-particle and multiparticle quantum batteries based on homodyne quantum feedback control. In the single-particle model, both the charger and the battery were modeled as two-level atomic systems. Quantum feedback control was applied to the charger during the charging process, enabling energy transfer through the coupling between the charger and the battery. Subsequently, we examined the effects of bosonic and fermionic thermal reservoirs on the performance of the single-particle quantum battery. The findings reveal the existence of optimal charging parameters for the system. Compared to bosonic thermal reservoirs, fermionic reservoirs at finite temperature facilitate energy storage and extraction more efficiently. Moreover, the battery charging efficiency $R$ improves with increasing temperature. In the multiparticle quantum battery model, the charger is still modeled as a two-level system, while the quantum battery consists of an ensemble of two-level systems. We further analyzed the charging process when the multiparticle quantum battery is collectively placed in either a bosonic or fermionic thermal reservoir. Additionally, we investigated the effects of the reservoir temperature $T$, the interparticle coupling strength $J$, and the number of particles $N$ on the performance of the multiparticle quantum battery. The results indicate that when the multiparticle quantum battery is placed in a bosonic thermal reservoir, increasing the temperature reduces both the energy density and the average ergotropy, consistent with the behavior observed in the single-particle quantum battery. Furthermore, expanding the size of the quantum battery in a bosonic thermal reservoir requires satisfying three conditions: a lower thermal reservoir temperature, weak inter-atomic interactions within the battery, and a low dissipation rate of the quantum battery. For the multiparticle quantum battery in a fermionic thermal reservoir, high temperatures still exhibit corresponding advantages in terms of the battery’s energy density and average ergotropy. Simultaneously, the conditions for constructing a multiparticle quantum battery are significantly relaxed. Even under strong interparticle interactions and higher battery dissipation rates, the multiparticle quantum battery still maintains relatively high-performance levels in both energy storage and energy extraction under high-temperature conditions.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant No. 12174048 and by Shenyang Ligong University through high-level talent research support funds under Grant No. 1010147001309.
Data availability
@@
Then setting $\dot{\rho}(t)=0$, we obtain the steady-state solutions
@@
$\mathcal{E}(t)$ $=$ $\frac{\omega_{0}}{2}(\sqrt{4|\rho_{12}(t)+\rho_{34}(t)|^{2}+[2(\rho_{11}(t)+\rho_{33}(t))-1]^{2}}+\{2[\rho_{11}(t)+\rho_{33}(t)]-1\}.$ (44)
According to Eqs. (43) and (44), the energy storage and the ergotropy of the quantum battery in the steady state can be determined. The expression for the ergotropy of the battery is more complex and will not be provided here, while the expression for energy storage is
@@
When the quantum battery is in contact with a fermionic thermal reservoir, the differential equations are similarly obtained by solving the corresponding master equation. By setting $\dot{\rho}(t)=0$, the steady-state solution is obtained as
@@
We set $\delta=f/\Gamma_{C}$ for simplicity. The other steady-state expansion coefficients are all zero, i.e. $\rho_{12}^{\infty}=\rho_{21}^{\infty}=\rho_{13}^{\infty}=\rho_{31}^{\infty}=\rho_{24}^{\infty}=\rho_{42}^{\infty}=\rho_{34}^{\infty}=\rho_{43}^{\infty}=0$. According to Eqs. (43) and (44), the stored energy and the ergotropy of the quantum battery in steady state can be determined. The expression for the ergotropy of the battery is more complex and will not be provided here, while the expression for energy storage is
@@
Appendix B Comparative analysis between the present work and Ref.  Song et al. (2024b)
In Ref.  Song et al. (2024b), the charger is modeled as a dissipative system embedded in either a bosonic or fermionic thermal reservoir, and powered by a driving laser field, while the quantum battery is isolated. For the case where the charger is embedded in a fermionic thermal reservoir with positive temperatures, we evaluate the quantum battery’s performance and conduct a comparative analysis with our proposed scheme. When the system reaches a steady state, the stored energy $E_{B}(\infty)$ of the battery and the ergotropy $\mathcal{E}(\infty)$ of the battery are,
$\mathcal{E}(\infty)$ $=$ $\frac{\omega_{0}}{2}\left(-1+2\sqrt{\frac{g^{2}[B+g^{2}(1-4n_{f})](1-2n_{f})^{2}\Gamma_{C}^{4}(2g^{2}+\Gamma_{C}^{2})^{2}}{\{4F^{4}g^{2}+2[A+2g^{4}(1-2n_{f})]\Gamma_{C}^{2}+[B+2g^{2}(1-2n_{f})]\Gamma_{C}^{4}\}^{2}}}\right.$ (53)
$\left.+\frac{4F^{4}g^{2}+2A\Gamma_{C}^{2}+B\Gamma_{C}^{4}}{4F^{4}g^{2}+2[A+2g^{4}(1-2n_{f})]\Gamma_{C}^{2}+[B+2g^{2}(1-2n_{f})]\Gamma_{C}^{4}}\right),$
Figure 22: Steady-state stored energy (solid line) and ergotropy (dashed line) of the quantum battery plotted as a function of the fermionic thermal reservoir temperature under optimal charging parameters $\Gamma_{C}=2g$ and $\delta=f/\Gamma_{C}=1$. The coupling strength between the charger and the battery is set to $g=0.01\omega_{0}$. The red curve represents the results from our proposed scheme. The corresponding results from Ref. Song et al. (2024b) are shown as the blue curve in the inset, with the parameter choices matching those in Fig. 5(a) of Ref. Song et al. (2024b), i.e.,$F=\Gamma_{C}=4g$.
Figure 23: The multiparticle quantum battery is placed in a bosonic thermal reservoir, with a strong dissipation rate of $\Gamma_{B}=0.5g$. The optimal energy density of the battery is plotted as a function of the number of quantum battery particles, under different reservoir temperatures and internal interactions between particles within the battery. The remaining information is provided in Fig. 14.
@@
Figure 25: The multiparticle quantum battery is in a bosonic thermal reservoir. The charging efficiency $R$ of the battery is shown as a function of the number of battery particles $N$. The remaining information is the same as in Fig. 16.
@@
respectively, where $A=F^{2}g^{2}[1-8(n_{f}-1)n_{f}]+F^{4}(1-2n_{f})^{2}+4g^{4}n_{f}$ and $B=F^{2}(1-2n_{f})^{2}+4g^{2}n_{f}$. Here $F$ is the amplitude of the driving field, $\Gamma_{C}$ is the dissipation rate of the charger, $g$ is the coupling strength between the charger and the battery, and $n_{f}$ is the average excitation number of the fermionic thermal reservoir.
@@
Figure 26: The multiparticle quantum battery is placed in a fermionic thermal reservoir. The optimal energy density of multiparticle quantum batteries is shown as a function of the particle number $N$. The only difference from Fig. 18 is that the dissipation rate of the battery is set to $\Gamma_{B}=0.5g$.
@@
To better compare the differences between our scheme and that in the Ref. Song et al. (2024b), Fig. 22 plots the steady-state stored energy of the quantum battery (red solid line) and its ergotropy (red dashed line) under our charging scheme in a fermionic thermal reservoir as a function of temperature. The corresponding results from Ref. Song et al. (2024b) are displayed in the inset, with the parameter choices matching those in Fig. 5(a) of Ref. Song et al. (2024b), i.e., $F=4g$ and $\Gamma_{C}=4g$. Here, the solid line represents the steady-state energy storage of the battery, and the dashed line corresponds to the maximum extractable energy (the ergotropy) of the battery. The results indicate that, in our scheme, both the energy storage and the maximum extractable energy of the battery increase with the temperature rise. In contrast, the results in Ref. Song et al. (2024b) show that, as the temperature increases, the maximum stored energy $E_{B}^{\rm optimal}(\infty)$ of the quantum battery tends to approach $0.5\omega_{0}$, and the maximum extractable energy $\mathcal{E}(\infty)$ gradually decreases to zero.
@@
Appendix C A strongly dissipative multiparticle quantum battery in a bosonic thermal reservoir
@@
Figure 27: In a fermionic thermal reservoir, the average ergotropy of the multiparticle quantum battery in the steady state is shown as a function of the particle number $N$. Except for $\Gamma_{B}=0.5g$, all other parameters are the same as in Fig. 19.
@@
Figure 28: The multiparticle quantum battery is coupled to a fermionic thermal reservoir. The charging efficiency $R$ of the battery is shown as a function of the number of battery particles $N$, where $\Gamma_{B}=0.5g$, and the other parameters are the same as those in Fig. 20.
@@
In this Appendix, we investigate the dependence of the energy density, the average ergotropy, and the charging efficiency $R$ of the multiparticle quantum battery on three key parameters, i.e., the thermal reservoir temperature, the number of particles $N$, and the interatomic interaction strength $J$, under strong dissipation conditions ($\Gamma_{B}=0.5g$) in a bosonic thermal reservoir. Figure 23 shows how the energy density of the multiparticle quantum battery varies with the particle number $N$ under different temperatures and interatomic interactions within the battery. Figures 23(a)-23(c) correspond to $J=0$, $0.1g$, and $g$, respectively, and the different curves represent different thermal reservoir temperatures. Figures 23(d)-23(f) correspond to the optimal dissipative rates of the charger required for Figs. 23(a)-23(c). Unlike the weak dissipation regime depicted in Fig. 14, the strong dissipation regime exhibits a universal decrease in energy density with increasing particle number $N$, independent of both the thermal reservoir temperature (ranging from low to high) and the strength of interparticle coupling within the battery (whether weak or strong). Consequently, constructing multiparticle quantum batteries in a strongly dissipative bosonic thermal reservoir is not advisable. However, it should be noted that for the number of particles $N>2$, a high-temperature environment can alleviate the reduction in energy density induced by increasing $N$. Figure 24 shows correspondingly the behavior of the average ergotropy of the multiparticle quantum battery. The result shows that in a high-temperature environment, an increase in the number of particles $N$ is beneficial for energy extraction. Meanwhile, as shown in Fig. 25, the charging efficiency $R$ of the battery increases with the number of particles under high-temperature conditions. In contrast, under low-temperature conditions, $R$ decreases as the number of particles $N$ increases.
@@
Appendix D A strongly dissipative multiparticle quantum battery in a fermionic thermal reservoir.
@@
Figure 26 shows the energy density of the battery as a function of the number of particles $N$ for the multiparticle quantum battery operating in a strongly dissipative fermionic thermal reservoir ($\Gamma_{B}=0.5g$). Figures 26(a)-26(c) correspond to $J=0$, $0.1g$, and $g$, respectively, with different curves representing various temperatures of the thermal reservoir. Figures 26(d)-26(f) show the corresponding optimal dissipative rates of the charger required for the cases in Figs. 26(a)-26(c), respectively. The results indicate that in a strong fermionic thermal reservoir, an increase in temperature enhances the energy density of the quantum battery. This means that higher temperatures facilitate energy storage in multiparticle quantum batteries. However, increasing the number of particles in the battery under low temperature conditions will have a negative impact on the energy density, regardless of whether the interatomic interaction strength is strong or weak. Conversely, under high-temperature conditions, as the number of battery particles increases, the battery will maintain a state with higher energy density, and the interatomic interaction strength $J$ has a relatively minor influence on its energy density. As shown in Fig. 27, the behavior of the average ergotropy closely resembles that of energy storage. Furthermore, Fig. 28 shows that the charging efficiency $R$ of the battery increases with the number of particles under high-temperature conditions, whereas under low-temperature conditions, $R$ decreases as $N$ increases.
@@
Overall, regardless of whether the multiparticle quantum battery operates under a low or high dissipation rate, its performance in a high-temperature fermionic thermal reservoir is significantly superior to that of a counterpart constructed in a bosonic thermal reservoir.